OpenStudy (anonymous):

How do you convert y-5=-2(x+1)^2 to standard and intercept modes?

OpenStudy (tkhunny):

Well, add 5 and that's usually Vertex Form. Then, multiply out and simplify the right side for Standard Form.

OpenStudy (anonymous):

Yeah, but the factoring is what is so tricky

OpenStudy (tkhunny):

Factoring? There is no need to factor. It is already in a factored form. Do you need to find zeros?

OpenStudy (anonymous):

yes, the form it currently holds is vertex. To get to intercept form, You must factor from standard

OpenStudy (tkhunny):

Factor or solve. Put it in Standard Form and see how easy it is.

OpenStudy (anonymous):

I did put it into standard form, and I got y= -2x^2-4x+3 remember the y-5 on the other side

OpenStudy (tkhunny):

y-5=-2(x+1)^2 y-5 = -2(x^2 + 2x + 1) y-5 = -2x^2 - 4x - 2 y = -2x^2 - 4x +3 Do you know the Discriminant? It will tell you that this cannot be factored. Just use the Quadratic Formula.

OpenStudy (anonymous):

So how am I supposed to answer the question for putting it in standard form? Also, I need to graph it and I'm just getting confused with what they expect here

OpenStudy (anonymous):

My own idiocy surprises me sometimes, it is able to be graphed.

OpenStudy (anonymous):

Thanks so much